Hi so in my problem, I have two shapes: a square and a rectangle. The square's dimensions are 5r and 5r. The rectangle's dimensions are 10r and 5r. For the square, I'm asked to find the area. Would the area for the square be 5r units squared or 25r units squared. And for the rectangle, would the answer be 50r units squared?
Answer by mgmoeab(37)

The sides of the square are 5r and 5r.
That means that the area of the square is A= S^2
Solution
if s= 5r
then A= s^2 ==== A =(5r)^2 = (5)^2 x (r)^2 = 25r^2
The sides of the rectangle are 10r and 5r.
That means that the area of the square is A= LxW
Solution
if L= 10r
W= 5r
then A= LxW ==== A =(10r)x (5r) = 50r^2
hope it helps!!